OpenStudy (anonymous):

Which of the following are the x and y intercepts of the equation 2x -2y = 3? A(2, 0), (0, 2) B.(1.5, 0), (0, -1.5) C.(2/3, 0), (0, -2/3) D.(3, 0), (0, -3)

OpenStudy (anonymous):

To find an x-intercept we substitute 0 for y in the equation and solve for x.

OpenStudy (anonymous):

So for an x-int, 2x-2y=3 becomes: 2x-2(0)=3 2x=3 x=3/2 = 1.5 So the x intercept is (1.5, 0)

OpenStudy (anonymous):

 Because the x-int occurs when y=0